OpenStudy (shamil98):

The population of deer in a certain national park can be approximated by the function P(x) = 150 (1.07)^x , where x is the number of years since 1995. In which year will the population reach 300?

ganeshie8 (ganeshie8):

300 = 150 (1.07)^x take log both sides ?

OpenStudy (shamil98):

2 = (1.07)^x log 2 / log 1.07 = x ?

ganeshie8 (ganeshie8):

yup! use calculator

ganeshie8 (ganeshie8):

answer should be x + 1995

OpenStudy (shamil98):

10.2448 = x

ganeshie8 (ganeshie8):

yes add 1995 to that

OpenStudy (shamil98):

oh, so 2005 then?

ganeshie8 (ganeshie8):

yes, in 2005 populatioon will reach 300